OpenStudy (anonymous):

A child is running on a moving sidewalk in an airport. When she runs against the sidewalk's motion, she travels at 40 ft in 8 seconds. When she runs with the sidewalk's motion, she travels at 132 ft in 12 seconds. What is the rate of the child on a still sidewalk and what is the rate of the moving sidewalk?

OpenStudy (anonymous):

Any type of explanation or answer will be greatly appreciated!

OpenStudy (anonymous):

Child's speed is \(c\), and sidewalk speed is \(s\).

OpenStudy (anonymous):

\[ c-s = \frac{40\;ft}{8\;s} \]And \[ c+s = \frac{132 \;ft}{12\;s} \]

OpenStudy (anonymous):

Do you think you can solve it now?